A 720 ml container holds a mixture of red, yellow and white paint in a 5 : 3 : 12 ratio. How much more red paint than yellow is present in the mixture?
Answer by Alan3354(69443)

A 720 ml container holds a mixture of red, yellow and white paint in a 5:3:12 ratio. How much more red paint than yellow is present in the mixture?
--------------
Assuming they're in that order, or respectively:
5+3+12 = 20
720/20 = 36
red:5*36 = 180 ml
ylw:3*36 = 108 ml
wht:12*36 = 432 ml
--------
red-ylw = 72 ml